Find the relative extremes of f?(x, y)=? ? x - x^2y - xy^2.
Answer #1

solution 11 from y) = x - 2² y - xy² The partial derivatives are fx (xuy) = 1- fy (x,y) = -x²-ray. f419,4)=0 1-2xy - y² |- 23